OpenStudy (anonymous):

Pump 1 can empty a pool in 9 hours while Pump 2 can empty the same pool in 12 hours. If the two pumps worked together, how long would it take them to empty the pool?

OpenStudy (anonymous):

Solve the following inequality: 2x + 6 ≤ 5x - 9

OpenStudy (anonymous):

Pump 1 does 1/9 per hour and Pump 2 does 1/12 per hour

OpenStudy (anonymous):

how many hours then together?

OpenStudy (anonymous):

Work out how much the two are doing per hour together and u can easily get your answer...

OpenStudy (anonymous):

about 5 hours about 5 hours and 10 minutes about 5 hours and 20 minutes about 5 hours and 30 minutes ?

OpenStudy (anonymous):

\[\frac{ 1pool }{ 9 hours }*T hours + \frac{ 1pool }{ 12 hours}*Thours = 1 pool\] t/9 + t/12 =1

OpenStudy (anonymous):

solve for t

OpenStudy (anonymous):

5 hours?

OpenStudy (anonymous):

you solved for t and got t=5?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

how do i get the t?

OpenStudy (anonymous):

multiply the whole equation by (9*12)

OpenStudy (anonymous):

108

OpenStudy (anonymous):

yes, multiply the whole equation by 108

OpenStudy (anonymous):

it will be easier for you if you leave 108 as 9*12 , but do it however you like...

OpenStudy (anonymous):

so the answerr is 5hr and 30mns

OpenStudy (anonymous):

what did you get for t?

OpenStudy (anonymous):

A plane traveling from Dallas to New York with a tailwind of 30 miles per hour takes 4.5 hours. Returning with the headwind, it takes the same plane 5 hours. What is the average speed of the plane in calm conditions? A plane traveling from Dallas to New York with a tailwind of 30 miles per hour takes 4.5 hours. Returning with the headwind, it takes the same plane 5 hours. What is the average speed of the plane in calm conditions?

OpenStudy (anonymous):

go right to the answer, no agony \[\frac{9\times 12}{9+12}\]

OpenStudy (anonymous):

so its 9

OpenStudy (anonymous):

i seriously doubt \[\frac{9\times 12}{9+12}=9\]

OpenStudy (anonymous):

i got 5.1428567143

OpenStudy (anonymous):

so the answer is 5hrs?